OpenStudy (anonymous):

does anyone know an algebraic way to solve ln(t)-t=ln(9.21), a better question is there an algebraic way, I think there has to be

OpenStudy (turingtest):

I don't think there is a simple algebraic way to do this. There are perhaps more advanced techniques oh satellite know I bet

OpenStudy (anonymous):

\[\ln(t)-t<0\] so there is no solution

OpenStudy (turingtest):

no real solution...

OpenStudy (anonymous):

you are not going to find that using algebra, i am almost certain

OpenStudy (anonymous):

any ideas how wolfram took it to imaginary plane

OpenStudy (anonymous):

ln(t)−t=2.22.....

OpenStudy (turingtest):

yeah can't proceed from there because of what sat pointed out

OpenStudy (anonymous):

what's the technique then to solve it going to imaginary numbers?

OpenStudy (turingtest):

It depends on the situation de moivre, complex analysis, etc... sometimes you can get imaginary numbers with just the quadratic formula

OpenStudy (anonymous):

you usually define \[\log(z)\] in the complex plane as \[\log(z)=\log(r)+i\theta\] but the function is not single values unless you specify \[0\leq \theta <2\pi\] or some other interval of length \[2\pi\] because the polar form of a complex number is not unique

OpenStudy (anonymous):

thanks for the help